Fill in the blank with a constant, so that the resulting expression can be factored as the product of two linear expressions:
2mn - 15mn + 6n + 3mn - 22mn + 7n + ___

Answer by ikleyn(53751)

.
Fill in the blank with a constant, so that the resulting expression can be factored
as the product of two linear expressions:
2mn - 15mn + 6n + 3mn - 22mn + 7n + ___
~~~~~~~~~~~~~~~~~~

2mn - 15mn + 6n + 3mn - 22mn + 7n = (2 - 15 + 3 - 22)mn + (6+7)n = -32mn + 13n = (-32m+13)*n.


The blank can be zero, and then  (-32+13)*n is the desired decomposition into the product 
of two linear expressions  (-32m+13) and n.